# Just ASR + LLM? A Study on Speech Large Language Models’ Ability to Identify and Understand Speaker in Spoken Dialogue

### v1

@@ Abstract
In recent years, we have observed a rapid advancement in speech language models (SpeechLLMs), catching up with humans’ listening and reasoning abilities. Remarkably, SpeechLLMs have demonstrated impressive spoken dialogue question-answering (SQA) performance in benchmarks like Gaokao, the English listening test of the college entrance exam in China, which seemingly requires understanding both the spoken content and voice characteristics of speakers in a conversation. However, after carefully examining Gaokao’s questions, we find the correct answers to many questions can be inferred from the conversation context alone without identifying the speaker asked in the question. Our evaluation of state-of-the-art models Qwen-Audio and WavLLM in both Gaokao and our proposed ”What Do You Like?” dataset shows a significantly higher accuracy in these context-based questions than in identity-critical questions, which can only be answered correctly with correct speaker identification. Our results and analysis suggest that when solving SQA, the current SpeechLLMs exhibit limited speaker awareness from the audio and behave similarly to an LLM reasoning from the conversation transcription without sound. We propose that our definitions and automated classification of context-based and identity-critical questions could offer a more accurate evaluation framework of SpeechLLMs in SQA tasks.
@@ 1 Introduction
Speech large language models (SpeechLLMs), trained with thousands or more hours of data for various questions, have demonstrated state-of-the-art performance in many speech tasks, including automatic speech recognition, speech translation, and emotion recognition. One of the tasks that SpeechLLMs are believed to solve confidently is spoken dialogue understanding. In spoken dialogue understanding, two or multiple speakers are talking to each other, and SpeechLLMs are tested on what is said and who said what. To answer these questions correctly, SpeechLLMs are expected to comprehend both the semantic content of the conversation and the acoustic characteristics of speakers. The high performance of SpeechLLMs in SQA benchmarks like Gaokao suggests that SpeechLLMs possess both of these capabilities.
Fig. 1: An illustration of identity-critical and context-based questions. (M) and (F) indicate male and female.
However, upon careful inspection, we found SpeechLLMs constantly made mistakes on some surprisingly simple questions. We hypothesize that SpeechLLMs are worse at differentiating speakers’ voices than commonly assumed. To prove our hypothesis, we single out speaker-related SQA questions that require identifying a particular speaker’s voice characteristics as Identity-Critical Questions (ICQs) and the remaining questions that ask for conversation content solvable without identifying any speaker as Context-Based Questions (CBQs). We evaluate the performance of state-of-the-art SpeechLLMs Qwen-Audio [1] and WavLLM [2], as well as a cascaded ASR + LLM system unaware of any speaker acoustics, on both types of questions in the popular Gaokao benchmark and our proposed benchmark “What Do You Like?”. For the latter, we control the speakers, the content of the conversations, and the questions asked to investigate the behaviors of SpeechLLMs better. Results show that both SpeechLLMs and the ASR + LLM system perform significantly worse on ICQs than CBQs in both benchmarks. Our analysis suggests that current SpeechLLMs perform similarly to ASR + LLM: they hack CBQs with the content of the conversation and fail on ICQs, unable to differentiate speakers’ characteristics such as gender.
We hope our observation can inspire a reconsideration of current SpeechLLMs’ SQA ability and promote future development of stronger SpeechLLMs and more comprehensive SQA benchmarks.
@@ 2.1 SpeechLLMs
Speech large language models (SpeechLLMs) have received much attention since the success of text-only language models. LTU-AS [3], built upon LTU [4], features a whisper encoder [5] and a Llama language model [6]. After finetuning with low-rank adaptation (LoRA; [7]) on a speech and audio question-answering dataset with 9.6M (audio, question, answer) tuples, the model can perceive various speech and audio events and shows satisfactory results on a wide range of speech and audio-related tasks like audio classification, speech recognition, and emotion recognition. Qwen-audio [1] uses a Whisper+Qwen [8] architecture and uses a whisper-style multi-task training paradigm. WavLLM [2] utilizes a whisper encoder and a WavLM [9] to encode semantic and acoustic information respectively. A two-stage curriculum training approach is applied to augment the generalization ability of the language model.
@@ 2.2 Speech and Audio Understanding Benchmarks
SD-Eval [10] is a multidimensional benchmark for dialogue understanding focusing on four specific aspects: emotion, accent, environment, and age. Evaluation on SD-Eval shows that SpeechLLMs achieve better performance than cascaded ASR+LLM across all four aspects, but with a significant gap below LLM receiving ground-truth transcripts and paralinguistic/environmental labels. AudioBench [11] is another recently proposed benchmark for speech and audio understanding. Cascaded ASR+LLM system outperforms SpeechLLMs for speech understanding tasks like ASR and speech question-answering, while SpeechLLMs excel when the task includes non-speech sound or needs paralinguistic information, like audio captioning and gender recognition. It’s worth noting that in AudioBench, speech question-answering is identified as a speech-intensive task where the cascaded system performs better than all the SpeechLLMs tested including WavLLM, Qwen-audio, and Salmonn [12]. Over-reliance on querying the text information for speech question-answering could be one of the major reasons for SpeechLLMs’ degraded performance on SQA samples that require paralinguistic information.
@@ 3 Identity-Critical and Context-Based SQA Questions
Based on our preliminary experiments with SpeechLLMs, we observed failure cases when SpeechLLMs are confused about “who said what.” We hypothesize that SpeechLLMs may not fully leverage the speaker “who” information in the audio to answer these questions since the “what” part is usually mentioned by another speaker but mis-selected by SpeechLLMs. To analyze the speaker factor alone and eliminate other factors, we propose a categorization based on whether the question can be answered correctly with or without correct speaker identification from the audio. In this section, we first discuss identity-critical questions and context-based questions in detail with examples. Then, we introduce an automatic method to classify questions into one of them.
@@ 3.1 Identity-Critical Question (ICQ)
An ICQ demands an accurate identification of the speaker referenced in the question to answer correctly. When multiple speakers’ preferences/actions/events are mentioned, the model must capture “who said what” and leverage the target speaker’s information and linguistic content to provide the correct answer. In the ICQ example in Figure 1, the question requires the model to both recognize and differentiate the speaker preferences (“fish” and “beef”) from the content by their voices. In ignorance of speaker information in the audio, SpeechLLMs usually randomly output an answer from the two candidates.information in the audio.
@@ 3.2 Context-Based Questions (CBQ)
A CBQ does not require to identify the referenced speaker in the question. This is because either (1) only one preference/event/action is provided and associated to a single speaker in the conversation, or (2) the preference/event/action of the two speakers are the same. Thus, the model only need to understand “what is said” in the conversation and the correct answer can be directly inferred purely from the text-based context without any speaker information. For instance, in the dialogue in the middle of Figure 1, only a single potential action is mentioned (e.g., the man wants to conquer his laziness), suggesting that the answer will only be “do exercise”; in the other CBQ example, both speakers are disappointed by the slow service.
3.3 Automatic Classification of ICQ and CBQ in Speaker-related
To automatically classify two-party SQA examples into ICQ and CBQ, we propose to leverage the current state-of-the-art language model (e.g., GPT-4) to assist the annotation since previous work has shown that GPT-4 achieves human-level performance on the simpler tasks [13, 14]. Assume we have the full transcript of an audio file and their associated speaker gender for each speaker turn. We edit the mentioned speaker gender in the question to an opposite gender and add an additional option “none.” After that, we request GPT-4 [15] to answer the edited question based on the transcript with speaker gender tags.
There are three cases GPT-4 may answer:
  • If GPT-4 chooses the other option mentioned in the transcript, we define the question as ICQ. For the ICQ example in Figure 1, the question will ask the man’s preference and the answer will be fish.
  • If GPT-4 answers “none,” we define the example as CBQ since only one preference/action/event is associated with a single speaker.
@@ 4 SQA Datasets
In this section, we present our analysis of CBQ and ICQ on Gaokao, an SQA dataset with real speech used to evaluate SpeechLLMs in [2].We find that a large portion of the speaker-related questions of this dataset is categorized as CBQ. In addition, we describe how we curate a synthetic SQA dataset “What Do You Like?” which is much simpler than Gaokao, allowing a more controlled study of SpeechLLM performance. The full Gaokao dataset is available at WavLLM’s github repository  [^1] https://github.com/microsoft/SpeechT5/tree/main/WavLLM and the data for our experiments can be accessed at  [^2] https://github.com/wjk0925/slt2024-speechllm-speaker-understanding.
@@ 4.1 Gaokao
Gaokao contains 2000 spoken dialogues between a woman and a man, each comes with a multiple-choice question with three options. All questions use “woman” or “man” when referring to a speaker, thus by searching these two keywords, we select 919 speaker-related questions in the dataset for analysis.
To classify these 919 SQA questions as ICQ or CBQ, we first obtain the full transcriptions of the dialogues using a Whisper [5] based speaker diarization + speech recognition system  [^3] https://github.com/MahmoudAshraf97/whisper-diarization and the inaSpeechSegmenter toolkit [16], which has gender prediction capability. The results are manually verified and corrected. Example dialogues are in Figure 1. We then follow the approaches described in section 3 and find 773 of the 919 speaker-related questions are CBQs and the rest are 146 ICQs.
@@ 4.2 What Do You Like?
Since questions in Gaokao requires varying level of reasoning and some dialogues contain potential gender hints such as names and titles (Mrs., Ms., Mr, …), we curated a synthetic SQA dataset “What Do You Like?” that follows the format of Gaokao (i.e. male-female dialogues, three-option multiple choice questions), but is much simplier to allow a more controllable analysis of SpeechLLMs’ performance.
Fig. 2: Sample What Do You Like?; Speaker genders are randomly decided during the TTS process
4.2.1 Spoken Dialogues
The dataset contains 1000 single-turn spoken dialogues between a male speaker and a female speaker discussing their favorite subjects under a certain topic. We first generate the text dialogue by randomly selecting a topic and two subjects, both from a predetermined list. We then fill in them into a template to make the text dialogue. Finally we use a text-to-speech model, StyleTTS-2 [17], to convert the text dialogue into waveform. For TTS voices, we randomly pick a pair (one male and one female) from 5 different male voices and 5 different female voices in VCTK [18] and the train split of LibriTTS [19], which StyleTTS-2 is trained on. An example is shown in Figure 2.
4.2.2 Questions and Options
We follow the Gaokao dataset and use multiple choice questions for “What Do You Like?”. For the question part, we simply ask a random speaker’s favorite subject, and use man/woman to address them based on their randomly assigned gender during the TTS process. Before generating the options, we set different labels for all the subjects in the predetermined list; the asked speaker’s favorite subject is labeled as Asked, the other speaker’s favorite subject is labeled as Mentioned, and other subjects are labeled as Unmentioned. Then, we generate four different options for the question
Options 1 include “Asked,” “Mentioned,” and a random “Unmentioned” in options.
Options 2 include “Asked” and two “Unmentioned” in options.
Options 3 include “Mentioned”, a random “Unmentioned,” and “None of the above” in options.
Options 4 include two “Unmentioned” and “None of the above” in options.
An example is shown in Figure 2. We can see although the question is ICQ if asked in a open-end way, options 2 and options 4 change it into a CBQ. This allows us to analyze SpeechLLMs’ performance under different questions categories. Moreover, the dialogues contain no obvious gender hints and the questions only require simple information extraction to solve, making it possible to focus on SpeechLLMs’ ability to identify speakers from voice differences.
@@ 5.1.1 SpeechLLMs
We focus on WavLLM and Qwen-Audio as they are reported as the top SpeechLLMs for SQA tasks in [2, 11]. Qwen-Audio often gives free-form responses to multiple-choice questions instead of one option. To obtain the final answer, we prompt GPT-4 with the multiple-choice question and the Qwen-Audio’s response, and have it output one of ‘A’, ‘B’, ‘C’, ‘X’, where ‘X’ means the response from Qwen-Audio does not correspond to any one of the choices A, B, and C.
5.1.2 ASR + LLM systems
@@ 5.2 Gaokao Results
[…]
We can see both WavLLM and Qwen-Audio exhibit significant poorer performance on ICQs compared with CBQs. The peformance gaps are similar to the two ASR + LLM cascaded systems, which have completely no access to the speaker acoustic information. Moreover, speaker diarization information only without gender label is not sufficient to bring much improvement, but with both speaker and gender information, the upper model scores higher on ICQs than CBQs, indicating a similar level of reasoning difficulty on the two splits. Our results show WavLLM and Qwen-Audio lack speaker identification ability required for ICQs but is capable of solving CBQs by inference with context.
@@ 5.3 Results on What Do You Like?
[…]
Fig. 3: Options 1 and Options 2
In Figure 3, we compare models’ performance on Options 1 (O1: Asked, Mentioned, Unmentioned) and Opitons 2 (O2: Asked, Unmentioned, Unmentioned). All SpeechLLMs and LLMs perform a random guess on O1 and reach a much higher accuracy on O2. As the only difference between O1 and O2 is the presence of the Mentioned option, this shows SpeechLLMs cannot utilize the voice information in the audio input to identify the asked speaker; instead they extract information from the dialogue content like LLMs. We note also note that Qwen-Audio and Llama3 choose the Unrelated option on O2 for a noticable portion of questions. This is likely because they are not trained to answer multiple-choice questions in the format of our data, and their responses are not always restricted by the provided options.
Fig. 4: Options 3 and Options 4
Figure 4 presents models’ performance on Options 3 (O3: Mentioned, Unmentioned, None of the above) and Options 4 (O4: Unmentioned, Unmentioned, None of the above). For simplicity, we will refer to None of the above as None. The correct answer in this case is None, and the difference between O3 and O4 is again the presence of Mentioned. We observe that on O3, all models have a strong intention to choose Mentioned than None, in contrast to the random guess behavior on O1. On O4, Qwen-Audio responses Unrelated for most of the questions, Llama3 reaches a nearly perfect performance, and the other two models show a much stronger intention to choose None. These are likely caused by the lack of None option in the training data of WavLLM and Qwen-Audio.

### v2

@@ Abstract
In recent years, we have observed a rapid advancement in speech language models (SpeechLLMs), catching up with humans’ listening and reasoning abilities. SpeechLLMs have demonstrated impressive spoken dialog question-answering (SQA) performance in benchmarks like Gaokao, the English listening test of the college entrance exam in China, which seemingly requires understanding both the spoken content and voice characteristics of speakers in a conversation. However, after carefully examining Gaokao’s questions, we find the correct answers to many questions can be inferred from the conversation transcript alone, i.e. without speaker segmentation and identification. Our evaluation of state-of-the-art models Qwen-Audio and WavLLM on both Gaokao and our proposed ”What Do You Like?” dataset shows a significantly higher accuracy in these context-based questions than in identity-critical questions, which can only be answered reliably with correct speaker identification. The results and analysis suggest that when solving SQA, the current SpeechLLMs exhibit limited speaker awareness from the audio and behave similarly to an LLM reasoning from the conversation transcription without sound. We propose that tasks focused on identity-critical questions could offer a more accurate evaluation framework of SpeechLLMs in SQA.
@@ 1 Introduction
Speech large language models (SpeechLLMs), trained with thousands or more hours of data for various questions, have demonstrated state-of-the-art performance in many speech tasks, including automatic speech recognition, speech translation, and emotion recognition. One of the tasks that SpeechLLMs are believed to solve reliably is spoken dialogue understanding. In spoken dialogue understanding, two or more speakers are talking to each other, and SpeechLLMs are tested on what is said and who said what. To answer these questions correctly, SpeechLLMs are expected to comprehend both the semantic content of the conversation and the acoustic characteristics of speakers. The high performance of SpeechLLMs in SQA benchmarks like Gaokao suggests that SpeechLLMs possess both of these capabilities.
Fig. 1: Examples of an identity-critical question (ICQ) and two context-based questions (CBQs). The male (M) and female (F) indicators in the transcript are only included in oracle experiments.
However, upon careful inspection, we found SpeechLLMs frequently made mistakes on some surprisingly simple questions. We hypothesize that SpeechLLMs are worse at differentiating speakers’ voices than commonly assumed. To test our hypothesis, we separate out speaker-related SQA questions that require identifying a particular speaker’s voice characteristics as Identity-Critical Questions (ICQs) and the remaining questions that ask for conversation content solvable without identifying any speaker as Context-Based Questions (CBQs). We evaluate the performance of state-of-the-art SpeechLLMs Qwen-Audio [1] and WavLLM [2], as well as cascaded ASR + LLM systems unaware of any speaker acoustics, on both types of questions in the popular Gaokao benchmark and our proposed benchmark “What Do You Like?”. For the latter, we control the speakers, the content of the conversations, and the questions asked to investigate the behaviors of SpeechLLMs better. Results show that both SpeechLLMs and the ASR + LLM system perform significantly worse on ICQs than CBQs in both benchmarks. Our analysis suggests that current SpeechLLMs perform similarly to ASR + LLM: they successfully answer CBQs with the content of the conversation and fail on ICQs, unable to differentiate speakers’ characteristics such as gender.
We hope our observations can inspire a reconsideration of current SpeechLLMs’ SQA ability and promote future development of stronger SpeechLLMs and more comprehensive SQA benchmarks.
@@ 2.1 SpeechLLMs
Speech large language models (SpeechLLMs) have received much attention since the success of text-only language models. LTU-AS [3], built upon LTU [4], features a Whisper encoder [5] and a LLaMA language model [6]. After finetuning with low-rank adaptation (LoRA; [7]) on a speech and audio question-answering dataset with 9.6M (audio, question, answer) tuples, the model can perceive various speech and audio events and shows satisfactory results on a wide range of speech and audio-related tasks like audio classification, speech recognition, and emotion recognition. Qwen-audio [1] uses a Whisper+Qwen [8] architecture and a Whisper-style multi-task training paradigm. WavLLM [2] utilizes a Whisper encoder and a WavLM [9] to encode semantic and acoustic information respectively. A two-stage curriculum training approach is applied to augment the generalization ability of the language model.
@@ 2.2 Speech and Audio Understanding Benchmarks
SD-Eval [10] is a multidimensional benchmark for dialogue understanding focusing on four specific aspects: emotion, accent, environment, and age. Evaluation on SD-Eval shows that SpeechLLMs achieve better performance than cascaded ASR+LLM systems across all four aspects, but with a significant gap below the LLM receiving ground-truth transcripts and paralinguistic/environmental labels. AudioBench [11] is another recently proposed benchmark for speech and audio understanding. Cascaded ASR+LLM systems outperform SpeechLLMs for speech understanding tasks like ASR and speech question-answering, while SpeechLLMs excel when the task includes non-speech sound or needs paralinguistic information, like audio captioning and gender recognition. It is worth noting that in AudioBench, speech question-answering is identified as a speech-intensive task where the cascaded system performs better than all the SpeechLLMs tested including WavLLM, Qwen-audio, and SALMONN [12]. Over-reliance on querying the text information for speech question-answering could be one of the major reasons for SpeechLLMs’ degraded performance on SQA samples that require paralinguistic information.
@@ 3 Identity-Critical and Context-Based SQA Questions
Based on our preliminary experiments with SpeechLLMs, we observed failure cases when SpeechLLMs are confused about “who said what.” We hypothesize that SpeechLLMs may not fully leverage the speaker “who” information in the audio to answer these questions, since the “what” part is usually mentioned by another speaker but mis-selected by SpeechLLMs. To analyze the speaker factor alone and eliminate other factors, we propose a categorization of questions based on whether they can be answered correctly with or without correct speaker identification from the audio. In this section, we describe identity-critical questions (ICQs) and context-based questions (CBQs), then introduce an automatic method to classify each question as an ICQ or a CBQ.
@@ 3.1 Identity-Critical Question (ICQ)
An ICQ demands an accurate identification of the speaker referenced in the question to answer correctly. When multiple speakers’ preferences/actions/events are mentioned, the model must capture “who said what” and leverage the target speaker’s information and linguistic content to provide the correct answer. In the ICQ example in Figure 1, the question requires the model to both recognize and differentiate the speaker preferences (“fish” and “beef”) from the content by their voices. In ignorance of speaker information in the audio, SpeechLLMs usually randomly output an answer from the two candidates mentioned in the audio.
@@ 3.2 Context-Based Questions (CBQ)
A CBQ does not require the identity of the speaker referenced in the question. This is because either: (1) only one preference/event/action is provided and associated to a single speaker in the conversation, or (2) the preference/event/action of the two speakers are the same. Thus, the model only needs to understand “what is said” in the conversation; the correct answer can be directly inferred purely from the text-based context. For instance, in the dialogue in the middle of Figure 1, only a single potential action is mentioned (e.g., the man wants to conquer his laziness), suggesting that the answer will only be “do exercise”. In the other CBQ example, both speakers are disappointed by the slow service.
3.3 Automatic Classification of ICQ and CBQ
To automatically classify two-party SQA examples as ICQ or CBQ, we leverage a state-of-the-art language model (GPT-4) to assist the annotation, since previous work has shown that GPT-4 achieves human-level performance on simple tasks [13, 14]. Given the full transcript of an audio file with the associated speaker gender for each speaker turn, we replace the mentioned speaker gender in the question with the opposite gender and add an additional answer option “None of the above.” Then, we request GPT-4 [15] to answer the edited question based on the transcript with speaker gender tags. The question is classified as follows:
  • If GPT-4 chooses the gender mentioned in the revised question, we define the question as ICQ. For the ICQ example in Figure 1, the question will ask the man’s preference and the answer will be fish.
  • If GPT-4 answers “None of the above,” we define the example as CBQ since only one preference/action/event is associated with a single speaker.
@@ 4 SQA Datasets
In this section, we present our analysis of CBQ and ICQ on Gaokao, an SQA dataset with real speech used to evaluate SpeechLLMs in [2]. We find that a large portion of the speaker-related questions of this dataset is categorized as CBQ. In addition, we describe a simple synthetic SQA dataset, What Do You Like?, which we curate to allow a more controlled study of SpeechLLM performance. The full Gaokao dataset is available at WavLLM’s github repository, [^1] https://github.com/microsoft/SpeechT5/tree/main/WavLLM and the data for our experiments can be accessed at  [^2] https://github.com/wjk0925/slt2024-speechllm-speaker-understanding.
@@ 4.1 Gaokao
Gaokao contains 2000 spoken dialogs between a woman and a man, each comes with a multiple-choice question with three options. All questions use “woman” or “man” when referring to a speaker, thus by searching these two keywords, we select 919 speaker-related questions in the dataset for analysis. Example dialogues are in Figure 1.
To classify these 919 SQA questions as ICQ or CBQ, we first obtain the full transcriptions of the dialogues using a Whisper-based speaker diarization + speech recognition system [5] [^3] https://github.com/MahmoudAshraf97/whisper-diarization and the inaSpeechSegmenter toolkit [16], which has gender classification capability. The diarization and gender classification results are manually verified and corrected. We then follow the approaches described in section 3 and identify 773 CBQs and 146 ICQs.
@@ 4.2 What Do You Like?
Since questions in Gaokao require varying levels of reasoning and some dialogues contain potential gender hints such as names and titles (Mrs., Ms., Mr, …), we curated a synthetic SQA dataset, What Do You Like? that follows the format of Gaokao (i.e. male-female dialogues, three-option multiple choice questions), but is much simplier to allow a more controllable analysis of SpeechLLMs’ performance.
Fig. 2: Example generation of a question and different answer options for What Do You Like?. The correct answer for each condition (C1:C4) is indicated by an asterisk.
The dataset contains 1000 single-turn spoken dialogues between a male speaker and a female speaker discussing their favorite subjects under a certain topic. As shown in Figure!2, we first generate the text dialogue by randomly selecting a topic and two subjects, both from a predetermined list of 3 topics (food, beverage, city) and 4 subjects for each topic. We then fill in them into one of 8 templates to make the text dialogue. Finally we use a text-to-speech model, StyleTTS-2 [17], to convert the text dialogue into waveform. For TTS voices, we randomly pick a pair (one male and one female) from 5 different male voices and 5 different female voices in VCTK [18] and the train split of LibriTTS [19], which StyleTTS-2 is trained on.
We follow the Gaokao dataset and use multiple choice questions with three options for “What Do You Like?”. For the question, we simply ask a random speaker’s favorite subject using one of two templates: “What is the man/woman’s favorite topic?” and “Which topic does the man/woman like the most?”. We use man/woman to address the target speaker based on their randomly assigned gender during the TTS process. We generate four different sets of answer options according to the following conditions:
C1: (Asked, Mentioned, Unmentioned)
C2: (Asked, Unmentioned, Unmentioned)
C3: (Mentioned, Unmentioned, None)
C4: (Unmentioned, Unmentioned, None)
where “Asked” is the favorite subject of the target speaker, “Mentioned” is the favorite subject of the other speaker, “Unmentioned” is another subject from the list, and “None” is the phrase “None of the above.”
Note that the question is ICQ if asked in a open-end way, but conditions 2 and 4 change it into a CBQ. The dialogues contain no obvious gender hints and the questions only require simple information extraction to solve, making it possible to focus on SpeechLLMs’ ability to identify speakers from voice differences under different answer conditions.
@@ 5.1.1 SpeechLLMs
We focus on WavLLM and Qwen-Audio as they are reported as the top SpeechLLMs for SQA tasks in [2, 11]. Qwen-Audio often gives free-form responses to multiple-choice questions instead of one option. To obtain the final answer, we prompt GPT-4 with the multiple-choice question and the Qwen-Audio’s response, and have it output one of ‘A’, ‘B’, ‘C’, or ‘X’, where ‘X’ means the response from Qwen-Audio does not correspond to any one of the choices A, B, or C.
5.1.2 Text-only systems
@@ 5.3 Results on What Do You Like?
The What Do You Like? dataset was designed to have four different answer set conditions to probe the behavior of SpeechLLM vs. text-only models. The conditions vary whether the “Asked” subject is in the answer set (if not, the correct answer is “None”) and whether the “Mentioned” subject is in the set. The analyses below show the distribution of the three possible answers plus the non-compliant answer type “X.” Only non-oracle models are included; the Llama3+SG peformance is nearly perfect for all four conditions.
Fig. 3: Number of times models choose each option for Condition 1 (left) and Condition 2 (right)
In Figure 3, we compare the models’ performance on Condition 1 (C1: Asked, Mentioned, Unmentioned) and Condition 2 (C2: Asked, Unmentioned, Unmentioned). As expected, all models perform at chance on C1. All have higher accuracy on C2, where there is no distracting Mentioned option. This indicates that SpeechLLMs are not using the voice information in the audio input to identify the asked speaker; instead they extract information from the dialogue content like LLMs. When there are two Unmentioned subjects in the answer set (C2), Qwen-Audio and Llama3 produce the X option fairly often. It may be that the Unmentioned answers are sufficiently out of context that the models do not recognize this as a multiple choice task.
Fig. 4: Number of times models choose each option for Condition 3 (left) and Condition 4 (right)
Figure 4 presents the models’ performance on Condition 3 (C3: Mentioned, Unmentioned, None) and Condition 4 (C4: Unmentioned, Unmentioned, None). The correct answer in these cases is None, and the difference between C3 and C4 is again the presence of Mentioned. We observe that on C3, the two WavLLM models virtually always chose the Mentioned option, and the other models frequently choose it. Only Llama3 gets a sizable fraction correct. On C4, Qwen-Audio responses X for most of the questions, Llama3 reaches nearly perfect performance, and the two WavLLM models show a much stronger tendency to choose the correct answer None. These findings provide further support for the hypothesis that the SpeechLLMs are not using speaker voice information.